OpenStudy (anonymous):

128 square root of e980

OpenStudy (anonymous):

If you solve for "e" it would be what times 980 has a square root of 128 answer is 16.7183673

OpenStudy (pawanyadav):

Is it e^980 or e(980)

OpenStudy (vats-sagar):

could u elaborate the problem

OpenStudy (agreene):

The way you have this written is a bit ambiguous. Do you perhaps mean: \[128\sqrt{e^{980}}\]

OpenStudy (mrnood):

sqrt(x) = x^0.5 use propoerties of exponents to simplify your expression

OpenStudy (mrnood):

OR do you mean \[\sqrt[128]{e ^{980}}\] in which case it is the same as \[(e ^{980})^{\frac{ 1 }{ 128 }}\] and you can still use teh same rules....